# From Text to Talk: Audio-Language Model Needs Non-Autoregressive Joint Training

### v3

@@ A.3.2 Absorbing Discrete Diffusion
Discrete diffusion models offer a non-autoregressive alternative for sequence generation. We focus on absorbing discrete diffusion (Austin et al., 2021; Ou et al., 2024), which involves a forward corruption process and a learned reverse denoising process.
Forward Process.
@@
Time-Independent Score and the Denoising Objective.
@@
$\underbrace{\frac{p_{t}(\ldots,\hat{x}^{i}{=}v,\ldots)}{p_{t}(\ldots,x^{i}=[\mathbf{M}],\ldots)}}_{\text{concrete score}}~=~\underbrace{\frac{e^{-\bar{\sigma}(t)}}{1-e^{-\bar{\sigma}(t)}}}_{\text{time scalar}}\;\cdot\;\underbrace{p_{0}\!\big(v~\big|~UM\big)}_{\text{clean conditional probability}}.$ (16)
where $UM$ denotes the set of unmasked (visible) tokens in the corrupted sequence. This decomposition is crucial because it decouples the time-dependent dynamics from the data distribution. It implies that the model $q_{\theta}$ does not need to learn a complex function of time $t$. Instead, its sole objective is to learn to approximate the clean conditional distribution $p_{0}(v|UM)$, which is a static, time-independent property of the data. The learning task is thus simplified to a denoising objective: given a corrupted sequence with some tokens masked, predict the original tokens for the masked positions based on the visible context.
Equivalence to Any-Order Autoregressive Modeling.
This denoising perspective reveals a profound connection to autoregressive modeling. A standard AR model learns to predict a token based on a fixed, causal context. The diffusion model, through its denoising objective, learns to predict a token given an arbitrary context of unmasked tokens. This ability to condition on any subset of the context is the defining feature of an Any-Order Autoregressive Model (AO-ARM).
@@
where $\pi$ is a random permutation of the token indices. Therefore, training an absorbing discrete diffusion model is equivalent to training a powerful ensemble of autoregressive models that can operate in any order. This inherent flexibility is what enables parallel, non-autoregressive generation at inference time and makes it a suitable choice for modeling source-dependent modalities like audio.
@@ A.4 Block-wise Masked Diffusion Generation For Audio Tokens
For NAR audio generation, we employ a block-wise denoising approach adapted from Nie et al. (2025). Unlike full-sequence diffusion, it processes audio in fixed-length blocks, balancing parallelism and controllability.
@@
Algorithm 1 Block-wise Masked Diffusion for Autoregressive Audio Generation
@@ A.5 Illustration of Training Loss and Attention Design
Figure 3(a) shows the training loss of our framework: AR loss is applied to text spans, while NAR loss—based on discrete diffusion—is used for audio spans. Although we employ discrete diffusion, our framework is extensible to other NAR generation methods. Figure 3(b) visualizes the attention pattern described in Section 3.4.
@@ A.7.1 Evaluation Tasks
We leverage URO-Bench Yan et al. (2025) for a more comprehensive evaluation of our proposed method against existing baselines. URO-Bench is specifically designed for audio-in, audio-out tasks, directly simulating real-world conversational scenarios. In this framework, the spoken outputs from each model are first transcribed into text using Whisper-large Radford et al. (2022), and the resulting transcripts are then evaluated for correctness, coherence, and task alignment. This evaluation employs a hybrid scoring framework comprising three components: (1) an LLM-as-a-judge (originally implemented via commercial LLM APIs) to assess semantic correctness and task alignment; (2) rule-based metrics for automatic Word Error Rate (WER) computation; and (3) a fine-tuned emotion-aware model to evaluate the appropriateness of affective expression in spoken responses. Together, these components ensure that model outputs are judged not merely on surface-level textual fidelity, but on semantic accuracy, transcription quality, and emotional coherence.
URO-Bench structures its evaluation across two difficulty levels: Basic and Pro. Each level comprises three distinct categories: Understanding tasks, Reasoning tasks, and Oral Conversation tasks. This hierarchical design facilitates a fine-grained assessment of model capabilities across increasing levels of linguistic and cognitive demand, covering both single-round and multi-round scenarios. In our experiments, we report results on the English subset of URO-Bench’s evaluation set. Due to limited access to the original evaluation APIs (Gemini Flash and GPT-4o-mini), we substitute the LLM-as-a-judge component with thinking model Qwen3-30B-A3B, while keeping all other components—including the Whisper-based ASR pipeline and rule-based scoring—identical to the original implementation.
Furthermore, the benchmark integrates a perceptual quality evaluation mechanism. We employ the strong UTMOS Saeki et al. (2022) for the UTMOS score and DNSMOS Dubey et al. (2024) for the NMOS score evaluation, enabling the joint assessment of content accuracy and acoustic quality. Importantly, none of the URO-Bench data was used during training or validation, ensuring an unbiased assessment of generalization.
@@
In addition to URO-Bench, we also evaluate our model with the Audio-QA tasks established by Kimi-Audio Ding et al. (2025). Previous evaluation framework in Kimi-Audio assess Audio-QA performance using the text portion of interleaved outputs, which overlooks the fact that the audio output of an end-to-end speech model more directly reflects its ability to generate natural and semantically faithful responses. To address this limitation, we evaluate Audio-QA directly on the audio outputs of our framework by first applying an ASR model to transcribe the generated audio into text, where Whisper-Large-v3 Radford et al. (2022) is used for English audio and Paraformer-zh for Chinese audio, with a comparison of ASR performance across different models provided in Table 5. The transcribed text is then combined with the original QA queries and the ground truth answers and passed to a large scale reasoning model, Qwen3-235B-A30B, which serves as an LLM-as-a-Judge model to determine whether the response semantically matches the reference and to provide either a correctness label or a graded score. We report the average accuracy or score on the benchmark, and this evaluation pipeline provides a more faithful assessment of our model’s audio-to-audio QA ability in realistic conversational scenarios where speech serves as the output modality.
@@
To assess the model’s capability in aligning speech with textual representations, we evaluate it on the ASR task, where the model generates text transcriptions from input audio and performance is measured using word error rate (WER). A lower WER indicates more accurate recognition, which reflects not only strong ASR ability but also effective cross modal consistency achieved by our hybrid AR-NAR modeling framework.
@@
To assess the model’s capacity to comprehend complex or acoustically challenging audios, we evaluate its audio captioning (AAC) performance on two established benchmarks: Clotho-v2 Drossos et al. (2020) and MACS Martín-Morató and Mesaros (2021). We input audio clips and generate corresponding textual captions. The quality of these captions is then evaluated using Qwen3-30B-A3B Yang et al. (2025a), guided by the evaluation prompt introduced in CLAIR-A Wu et al. (2024), which emphasizes semantic relevance, completeness, and naturalness. The judge assigns a score on a 0–100 scale, where higher scores indicate better caption quality.
@@ A.8 Baselines
  • Moshi Défossez et al. (2024): It unifies streaming speech and text understanding within a single autoregressive framework, aligning acoustic and linguistic representations for low-latency real-time dialogue and robust multimodal instruction following.
  • SpeechGPT Zhang et al. (2023): It incorporates discrete speech tokens into a single language model and follows a three-stage training pipeline to enable unified speech–text understanding and cross-modal instruction following within one framework.
@@
  • LLaMA-Omni Fang et al. (2025): It extends a unified language model to support real-time speech understanding and generation by integrating low-latency audio streaming, codec-based tokenization, and tightly aligned speech–text representations for seamless multimodal interaction.
@@
  • Mini-Omni Xie and Wu (2024a): It enables real-time speech interaction by generating text and audio tokens in parallel within one model, using text-instructed parallel decoding and a lightweight training pipeline to preserve the base model’s reasoning ability.
  • SLAM-Omni Chen et al. (2025): It enables end-to-end spoken dialogue by modeling text and semantic audio tokens in parallel within a single model, supporting zero-shot timbre control and low-latency voice interaction through single-stage training.
@@ A.9 Training Details
We train our model using the AdamW optimizer with a global batch size of 2048, a learning rate of $2e^{-5}$, and a weight decay factor of $1e^{-2}$. The learning rate follows a cosine decay schedule with a linear warmup ratio of $0.01$. Training incorporates three stochastic strategies: (1) batchwise AR & NAR objective mixing with probability 0.3; (2) prefix preservation masking with ratio 0.3; (3) stochastic span truncation with probability 0.5. During inference, the model alternates between AR text decoding and NAR diffusion-based audio generation, where text decoding uses nucleus sampling with $k=10$ and $p=0.95$, and audio spans are generated with 200 diffusion steps, a block length of 32 tokens, and a total diffusion span length of 640 tokens under classifier-free guidance with scale 0.1. Since different training strategies lead to varying convergence speeds, reported results are based on checkpoints where training loss has converged. All experiments are conducted on 4 nodes with 8 NVIDIA A100 GPUs per node using the DeepSpeed runtime.
@@ A.9.1 Data Format of Training Data
To enable unified training across diverse tasks, we transform all datasets into a consistent input–output format. On the one hand, this standardization allows the model to seamlessly integrate heterogeneous modalities such as speech, text, and interleaved audio–text sequences. On the other hand, a unified design is essential for supporting our training strategies, including batchwise AR & NAR objective mixing, prefix preservation masking, and stochastic span truncation. These strategies rely on a shared representation to operate across modalities in a consistent way. For clarity, we provide representative examples of the adopted data formats as follow, covering ASR, TTS, audio chat, text chat, AAC, SEC, ASC, and interleaved text–audio data.
@@
Figure 7: Example of text chat data format.
@@
A.10 Usage of LLM
In this paper, the LLM is employed solely for text refinement—correcting typos, fixing spelling errors, and enhancing readability. It is not used for generating research ideas, producing results, or creating original content.

### v4

@@ A.3.2 Absorbing Discrete Diffusion
Discrete diffusion models offer a NAR alternative for sequence generation. We focus on absorbing discrete diffusion (Austin et al., 2021; Ou et al., 2024), which involves a forward corruption process and a learned reverse denoising process.
Forward Process
@@
Time-Independent Score and the Denoising Objective
@@
$\underbrace{\frac{p_{t}(\ldots,\hat{x}^{i}{=}v,\ldots)}{p_{t}(\ldots,x^{i}=[\mathbf{M}],\ldots)}}_{\text{concrete score}}~=~\underbrace{\frac{e^{-\bar{\sigma}(t)}}{1-e^{-\bar{\sigma}(t)}}}_{\text{time scalar}}\;\cdot\;\underbrace{p_{0}\!\big(v~\big|~x_{\mathrm{vis}}\big)}_{\text{clean conditional probability}}.$ (16)
where $x_{\mathrm{vis}}$ denotes the set of unmasked (visible) tokens in the corrupted sequence. This decomposition is crucial because it decouples the time-dependent dynamics from the data distribution. It implies that the model $q_{\theta}$ does not need to learn a complex function of time $t$. Instead, its sole objective is to learn to approximate the clean conditional distribution $p_{0}(v|x_{\mathrm{vis}})$, which is a static, time-independent property of the data. The learning task is thus simplified to a denoising objective: given a corrupted sequence with some tokens masked, predict the original tokens for the masked positions based on the visible context.
Equivalence to Any-Order AR Modeling
This denoising perspective reveals a profound connection to AR modeling. A standard AR model learns to predict a token based on a fixed, causal context. The diffusion model, through its denoising objective, learns to predict a token given an arbitrary context of unmasked tokens. This ability to condition on any subset of the context is the defining feature of an any-order AR model (AO-ARM).
@@
where $\pi$ is a random permutation of the token indices. Therefore, training an absorbing discrete diffusion model is equivalent to training a powerful ensemble of AR models that can operate in any order. This inherent flexibility is what enables parallel, NAR generation at inference time and makes it a suitable choice for modeling source-dependent modalities like audio.
@@ A.4 Block-wise Masked Diffusion Generation For Audio Tokens
For NAR audio generation, we employ a block-wise denoising approach adapted from (Nie et al., 2025). Unlike full-sequence diffusion, it processes audio in fixed-length blocks, balancing parallelism and controllability.
@@
Algorithm 1 Block-wise Masked Diffusion for NAR Audio Generation
@@ A.5 Illustration of Training Loss and Attention Design
Figure 3(a) shows the training loss of our framework: AR loss is applied to text spans, while NAR loss (based on discrete diffusion) is used for audio spans. Although we employ discrete diffusion, our framework is extensible to other NAR generation methods. Figure 3(b) visualizes the attention pattern described in Section 3.4.
@@ A.7.1 Evaluation Tasks
We leverage URO-Bench for a more comprehensive evaluation of our proposed method against existing baselines (Yan et al., 2025). URO-Bench is specifically designed for audio-in, audio-out tasks, directly simulating real-world conversational scenarios. In this framework, the spoken outputs from each model are first transcribed into text using Whisper-large-v3 (Radford et al., 2023), and the resulting transcripts are then evaluated for correctness, coherence, and task alignment. This evaluation employs a hybrid scoring framework comprising three components: (1) an LLM-as-a-judge (originally implemented via commercial LLM APIs) to assess semantic correctness and task alignment; (2) rule-based metrics for automatic WER computation; and (3) a fine-tuned emotion-aware model to evaluate the appropriateness of affective expression in spoken responses. Together, these components ensure that model outputs are judged not merely on surface-level textual fidelity, but on semantic accuracy, transcription quality, and emotional coherence.
URO-Bench structures its evaluation across two difficulty levels: Basic and pro. Each level comprises three distinct categories: Understanding tasks, Reasoning tasks, and Oral Conversation tasks. This hierarchical design facilitates a fine-grained assessment of model capabilities across increasing levels of linguistic and cognitive demand, covering both single-round and multi-round scenarios. In our experiments, we report results on the English subset of URO-Bench’s evaluation set. Due to limited access to the original evaluation APIs (Gemini Flash and GPT-4o-mini), we substitute the LLM-as-a-judge component with thinking model Qwen3-30B-A3B, while keeping all other components, including the Whisper-based ASR pipeline and rule-based scoring, identical to the original implementation.
Furthermore, the benchmark integrates a perceptual quality evaluation mechanism. We employ the strong UTMOS for the UTMOS score and DNSMOS for the NMOS score evaluation, enabling the joint assessment of content accuracy and acoustic quality (Saeki et al., 2022; Dubey et al., 2024). Importantly, none of the URO-Bench data was used during training or validation, ensuring an unbiased assessment of generalization.
@@
In addition to URO-Bench, we also evaluate our model with the Audio-QA tasks established by Kimi-Audio (Ding et al., 2025). Previous evaluation framework in Kimi-Audio assesses Audio-QA performance using the text portion of interleaved outputs, which overlooks the fact that the audio output of an end-to-end speech model more directly reflects its ability to generate natural and semantically faithful responses. To address this limitation, we evaluate Audio-QA directly on the audio outputs of our framework by first applying an ASR model to transcribe the generated audio into text, where Whisper-large-v3 is used for English audio and Paraformer-zh for Chinese audio, with a comparison of ASR performance across different models provided in Table 5. The transcribed text is then combined with the original QA queries and the ground truth answers and passed to a large scale reasoning model, Qwen3-30B-A3B, which serves as an LLM-as-a-Judge model to determine whether the response semantically matches the reference and to provide either a correctness label or a graded score. We report the average accuracy or score on the benchmark, and this evaluation pipeline provides a more faithful assessment of our model’s audio-to-audio QA ability in realistic conversational scenarios where speech serves as the output modality.
@@
To assess the model’s capability in aligning speech with textual representations, we evaluate it on the ASR task, where the model generates text transcriptions from input audio and performance is measured using WER. A lower WER indicates more accurate recognition, which reflects not only strong ASR ability but also effective cross modal consistency achieved by our hybrid AR-NAR modeling framework.
@@
To assess the model’s capacity to comprehend complex or acoustically challenging audios, we evaluate its audio captioning (AAC) performance on two established benchmarks: Clotho-v2 (Drossos et al., 2020) and MACS (Martín-Morató and Mesaros, 2021). We input audio clips and generate corresponding textual captions. The quality of these captions is then evaluated using Qwen3-30B-A3B (Yang et al., 2025a), guided by the evaluation prompt introduced in CLAIR-A (Wu et al., 2025), which emphasizes semantic relevance, completeness, and naturalness. The judge assigns a score on a 0-100 scale, where higher scores indicate better caption quality.
@@ A.8 Baselines
  • Moshi Défossez et al. (2024): It unifies streaming speech and text understanding within a single AR framework, aligning acoustic and linguistic representations for low-latency real-time dialogue and robust multimodal instruction following.
  • SpeechGPT Zhang et al. (2023): It incorporates discrete speech tokens into a single language model and follows a three-stage training pipeline to enable unified speech-text understanding and cross-modal instruction following within one framework.
@@
  • LLaMA-Omni Fang et al. (2025): It extends a unified language model to support real-time speech understanding and generation by integrating low-latency audio streaming, codec-based tokenization, and tightly aligned speech-text representations for seamless multimodal interaction.
@@
  • Mini-Omni Xie and Wu (2024): It enables real-time speech interaction by generating text and audio tokens in parallel within one model, using text-instructed parallel decoding and a lightweight training pipeline to preserve the base model’s reasoning ability.
  • SLAM-Omni Chen et al. (2025a): It enables end-to-end spoken dialogue by modeling text and semantic audio tokens in parallel within a single model, supporting zero-shot timbre control and low-latency voice interaction through single-stage training.
@@ A.9 Training Details
We train our model using the AdamW optimizer with a global batch size of 2048, a learning rate of $2e^{-5}$, and a weight decay factor of $1e^{-2}$. The learning rate follows a cosine decay schedule with a linear warmup ratio of $0.01$. Training incorporates three stochastic strategies: (1) BANOM with probability 0.3; (2) PPM with ratio 0.3; (3) SST with probability 0.5. During inference, the model alternates between AR text decoding and NAR diffusion-based audio generation, where text decoding uses nucleus sampling with $k=10$ and $p=0.95$, and audio spans are generated with 200 diffusion steps, a block length of 32 tokens, and a total diffusion span length of 640 tokens under classifier-free guidance with scale 0.1. Since different training strategies lead to varying convergence speeds, reported results are based on checkpoints where training loss has converged. All experiments are conducted on 4 nodes with 8 NVIDIA A100 GPUs per node using the DeepSpeed runtime.
@@ A.9.1 Data Format of Training Data
To enable unified training across diverse tasks, we transform all datasets into a consistent input-output format. On the one hand, this standardization allows the model to seamlessly integrate heterogeneous modalities such as speech, text, and interleaved audio-text sequences. On the other hand, a unified design is essential for supporting our training strategies, including batchwise AR & NAR objective mixing, prefix preservation masking, and stochastic span truncation. These strategies rely on a shared representation to operate across modalities in a consistent way. For clarity, we provide representative examples of the adopted data formats as follows, covering ASR, TTS, audio chat, text chat, AAC, SEC, ASC, and interleaved text-audio data, as shown in Figures 4 - 9.
A.10 Usage of LLM
In this paper, the LLM is employed solely for text refinement, including correcting typos, fixing spelling errors, and enhancing readability. It is not used for generating research ideas, producing results, or creating original content.
@@ A.10 Usage of LLM
Figure 9: Example of text chat data format.